OpenStudy (anonymous):

Simplify : 15m^3n^8/3mn^4 a. 5m^2n^4 b. 5m^3n^2 c. 5m^4n^12

OpenStudy (callisto):

\[\frac{15m^3n^8}{3mn^4}\]\[=\frac{15}{3}m^{3-1}n^{8-4}\]\[=...?\]

OpenStudy (anonymous):

5m2n4. Thanks @Callisto !

OpenStudy (callisto):

Welcome :)